Factorise. (x^2 –2xy+y^2 )–z^2 ?????????????????????

Factorise.
 (x^2 –2xy+y^2 )–z^2 ?????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

= (x–y)^2–z^2
Using Identity: (x-y)^2= x^2-2xy+y^2
= {(x–y)–z}{(x–y)+z}
= (x–y–z)(x–y+z)
Using Identity: x^2-y^2= (x+y)(x-y)

Thanks